# Mixture of Reasonings: Teach Large Language Models to Reason with Adaptive Strategies

### v1

@@ Abstract.
Large Language Models, Mixture of Reasoning, Natural Language Processing
copyright: acmlicensedjournalyear: 2025conference: Proceedings of the 48th International ACM ICMR Conference on Research and Development in Information Retrieval; June 30–July 3, 2025; Chicago, IL, USA.booktitle: Proceedings of the 48th International ACM ICMR Conference on Research and Development in Information Retrieval (ICMR ’25), June 30–July 3, 2025, Chicago, IL, USAisbn: 979-8-4007-1592-1/25/07ccs: Computing methodologies Natural language processing
1. Introduction
@@
Figure 1. Overview of our proposed MoR framework. The MoR framework can be divided into two stages: (1) Thought Generation. As shown in step 1, this involves generating a large number of reasoning chain templates using GPT. (2) SFT Dataset Construction. As depicted in steps 2, 3, and 4, this includes selecting optimal reasoning chains, creating prompts, and filtering for correct responses.
2. Related Work
@@ 3. Method
3.1. Thought Generation
For small parameter models, due to insufficient embedded knowledge and limited reasoning capabilities, simply instructing them with ”Let’s think step by step” does not effectively stimulate the model’s capabilities.
@@ 3.2. SFT Dataset Construction
Table 1. Performance on reasoning tasks. We selected Qwen2.5-7B-instruct as the baseline model. We train the baseline model using our MoR approach by varying the number of reasoning chain templates. Additionally, to evaluate the effectiveness of MoR, we extend the test set from 50 to 200 instances, specifically comparing the baseline model with $MoR_{150}$. The best results for each setting are bolded.
Figure 2. Case study comparing the baseline model and $MoR_{150}$ using CoT prompts. The Qwen2.5-7B-instruct model follows the ”Let’s think step by step.” approach but ultimately produces incorrect answers. In contrast, the $MoR_{150}$ model adopts the MoR reasoning method, analyzing problems logically and ultimately arriving at the correct answer.
4. Experiment
4.1. Setup
Datasets.
In the experiment, we selected five reasoning datasets, with 50 samples randomly chosen from each dataset for testing. For BigTom, we selected 20 samples across four different ”belief settings,” totaling 80 samples. The SFT dataset construction used the GPT-4o-2024-08-06 version of GPT, as mentioned in §3.
@@ 4.1. Setup
Model. We selected the Qwen2.5-7B-Instruct (Qwen et al., 2025) model as the baseline. The models fine-tuned on different numbers of X-chain of thought datasets are used as our comparison models, denoted as $MoR_{i},wherei=50,150,300,500$. We believe that after training, the model has acquired MoR capabilities, so simply using the prompt ”Let’s think step by step.” is sufficient to elicit the model’s multi-step reasoning ability. We refer to this prompting strategy as the CoT prompt. For comparison, we also provide a setting where the model is directly instructed to answer the question without any special prompt which is called the IO prompt.
4.2. Result

### v2

@@ Abstract
Mixture of Reasonings: Teach Large Language Models to Reason with Adaptive Strategies
Tao Xiong1, †, Xavier Hu2, †, Wenyan Fan2, Shengyu Zhang2, ‡ 1Dalian University of Technology, 2Zhejiang University Correspondence: 030130x@mail.dlut.edu.cn, sy_zhang@zju.edu.cn
footnotetext: Both authors contributed equally to this research.
Corresponding Author
1 Introduction
@@
Figure 1: Overview of our proposed MoR framework. The MoR framework can be divided into two stages: (1) Thought Generation. As shown in step 1, this involves generating a large number of reasoning chain templates using GPT. (2) SFT Dataset Construction. As depicted in steps 2, 3, and 4, this includes selecting optimal reasoning chains, creating prompts, and filtering for correct responses.
2 Related Work
@@ 3 Method
3.1 Thought Generation
For small parameter models, due to insufficient embedded knowledge and limited reasoning capabilities, simply instructing them with "Let’s think step by step" does not effectively stimulate the model’s capabilities.
@@ 3.2 SFT Dataset Construction
Table 1: Performance on reasoning tasks. We selected Qwen2.5-7B-instruct as the baseline model. We train the baseline model using our MoR approach by varying the number of reasoning chain templates. Additionally, to evaluate the effectiveness of MoR, we extend the test set from 50 to 200 instances, specifically comparing the baseline model with $MoR_{150}$. The best results for each setting are bolded.
Figure 2: Case study comparing the baseline model and $MoR_{150}$ using CoT prompts. The Qwen2.5-7B-instruct model follows the "Let’s think step by step." approach but ultimately produces incorrect answers. In contrast, the $MoR_{150}$ model adopts the MoR reasoning method, analyzing problems logically and ultimately arriving at the correct answer.
4 Experiment
4.1 Setup
Datasets. In the experiment, we selected five reasoning datasets, with 50 samples randomly chosen from each dataset for testing. For BigTom, we selected 20 samples across four different "belief settings," totaling 80 samples. The SFT dataset construction used the GPT-4o-2024-08-06 version of GPT, as mentioned in §3.
@@ 4.1 Setup
Model. We selected the Qwen2.5-7B-Instruct (Qwen et al., 2025) model as the baseline. The models fine-tuned on different numbers of X-chain of thought datasets are used as our comparison models, denoted as $MoR_{i},wherei=50,150,300,500$. We believe that after training, the model has acquired MoR capabilities, so simply using the prompt "Let’s think step by step." is sufficient to elicit the model’s multi-step reasoning ability. We refer to this prompting strategy as the CoT prompt. For comparison, we also provide a setting where the model is directly instructed to answer the question without any special prompt which is called the IO prompt.
4.2 Result